# Quattro: Transformer-Accelerated Iterative Linear Quadratic Regulator Framework for Fast Trajectory Optimization

### v1

@@ Abstract
Real-time optimal control remains a fundamental challenge in robotics, especially for nonlinear systems with stringent performance requirements. As one of the representative trajectory optimization algorithms, the iterative Linear Quadratic Regulator (iLQR) faces limitations due to their inherently sequential computational nature, which restricts the efficiency and applicability of real-time control for robotic systems. While existing parallel implementations aim to overcome the above limitations, they typically demand additional computational iterations and high-performance hardware, leading to only modest practical improvements. In this paper, we introduce Quattro, a transformer-accelerated iLQR framework employing an algorithm-hardware co-design strategy to predict intermediate feedback and feedforward matrices. It facilitates effective parallel computations on resource-constrained devices without sacrificing accuracy. Experiments on cart-pole and quadrotor systems show an algorithm-level acceleration of up to 5.3$\times$ and 27$\times$ per iteration, respectively. When integrated into a Model Predictive Control (MPC) framework, Quattro achieves overall speedups of 2.8$\times$ for the cart-pole and 17.8$\times$ for the quadrotor compared to the one that applies traditional iLQR. Transformer inference is deployed on FPGA to maximize performance, achieving up to 27.3$\times$ speedup over commonly used computing devices, with around 2 to 4$\times$ power reduction and acceptable hardware overhead.
@@ IV-B Transformer Model and Accelerator Design
Clock Frequency 400 MHz (Quadrotor) 100 MHz (Cart-Pole)
@@ IV-D Computation Speedup and Accelerator Performance
Compared to another accelerated iLQR algorithm in [12], which typically requires additional iterations to converge, our method achieves a comparable number of iterations as vanilla iLQR to reach the optimal solution, thereby demonstrating significant overall time savings. As depicted in the upper plot of Figure 5, the total simulation time for the cart-pole system is reduced from 10.5 s to 3.7 s when using 5 iLQR steps and 25 Transformer-assisted steps, resulting in a 2.8$\times$ speedup. This aligns closely with the observed time savings illustrated in Figure 4 (upper). For the quadrotor control task with 500 control steps, we observed a more effective improvement of computation time from 237 s (pure iLQR computation) to 13.3 s (1 step of iLQR computation and 49 steps Transformer prediction) and the latter is 17.8$\times$ faster. We compare our framework with the state-of-the-art solvers OSQP, ECOS, and SCS in [44] for solving the quadrotor OCP, and we observed speedups of 2.49$\times$, 3.63$\times$, and 2.4$\times$, respectively.
@@
Table II shows that our accelerator achieves the expected performance in Transformer inference. For the cart-pole control scenario, it is 1.47$\times$ faster than the Apple M4 Pro CPU, 2.4$\times$ faster than the NVIDIA RTX4070, and 27.3$\times$ faster than the Cortex-A72 in the Raspberry Pi. In the quadrotor control scenario, our accelerator provides speedups of 1.6$\times$ over the M4 Pro CPU and 18.7$\times$ over the Raspberry Pi. It is 0.2 ms (0.9$\times$) slower than the RTX4070 on the quadrotor problem, but the measured power consumption (8.11 W for cart-pole and 3.46 W for quadrotor) is significantly lower than the 13 W drawn by the GPU under both tasks.
Hardware overhead remains acceptable, particularly for the Quadrotor scenario, where less than half of the Kintex FPGA resources are used. Under the same 20nm process, the estimated silicon area is significantly smaller than that of typical CPUs or GPUs.
TABLE II: Accelerator Performance and Resource Utilization (Cart-Pole vs. Quadrotor)
@@
Latency (ms) 0.68 1.95
Speed-Up (vs. 10-core M4 Pro CPU) 1.47$\times$ 1.6$\times$
Speed-Up (vs. NVIDIA RTX4070) 2.4$\times$ 0.9$\times$
Speed-Up (vs. Quad-core Cortex-A72) 27.3$\times$ 18.74$\times$
@@
BRAM Usage (%) 49 48
DSP Usage (%) 89 31
FF Usage (%) 18 8
LUT Usage (%) 54 15
@@ V CONCLUSION
In this paper, we presented Quattro, a Transformer-accelerated iLQR framework that addresses the sequential bottleneck in traditional iLQR algorithms through parallel inference of control gains. By predicting intermediate computational variables, Quattro maintains the structure of optimal control while significantly reducing computation time. Our experiments on cart-pole and quadrotor systems demonstrate that Quattro achieves up to 27$\times$ per-iteration speedup and 17.8$\times$ end-to-end acceleration within an MPC framework. FPGA-based deployment further shows 18–27$\times$ speedups over edge CPUs and up to 2.4 $\times$ than GPU with lower power consumption and acceptable hardware overhead. While larger Transformer models can yield even higher prediction accuracy (sometimes surpassing the original iLQR), they introduce higher latency and hardware overhead. Therefore, we select a lightweight configuration (dmodel=128, nhead=4) that balances accuracy, efficiency, and employability. These results confirm the potential of Transformer-based acceleration for real-time optimal control. Future work includes extending Quattro to more complex robotic systems, investigating adaptive prompt tuning, and exploring integration with learning-based control for greater robustness in dynamic environments.

### v2

@@ Abstract
Real-time optimal control remains a fundamental challenge in robotics, especially for nonlinear systems with stringent performance requirements. As one of the representative trajectory optimization algorithms, the iterative Linear Quadratic Regulator (iLQR) faces limitations due to their inherently sequential computational nature, which restricts the efficiency and applicability of real-time control for robotic systems. While existing parallel implementations aim to overcome the above limitations, they typically demand additional computational iterations and high-performance hardware, leading to only modest practical improvements. In this paper, we introduce Quattro, a transformer-accelerated iLQR framework employing an algorithm-hardware co-design strategy to predict intermediate feedback and feedforward matrices. It facilitates effective parallel computations on resource-constrained devices without sacrificing accuracy. Experiments on cart-pole and quadrotor systems show an algorithm-level acceleration of up to 5.3$\times$ and 27$\times$ per iteration, respectively. When integrated into a Model Predictive Control (MPC) framework, Quattro achieves overall speedups of 2.8$\times$ for the cart-pole and 17.8$\times$ for the quadrotor compared to the one that applies traditional iLQR. Transformer inference is deployed on FPGA to maximize performance, achieving further up to 20.8$\times$ speedup over prevalent embedded CPUs with over 11$\times$ power reduction than GPU and low hardware resource overhead.
@@ IV-B Transformer Model and Accelerator Design
Clock Frequency 80 MHz (Cart-Pole) 200 MHz (Quadrotor)
@@ IV-D Computation Speedup and Accelerator Performance
Compared to another accelerated iLQR algorithm in [12], which typically requires additional iterations to converge, our method achieves a comparable number of iterations as vanilla iLQR to reach the optimal solution, thereby demonstrating significant overall time savings. As depicted in the upper plot of Figure 5, the total simulation time for the cart-pole system is reduced from 10.5 s to 3.7 s when using 5 iLQR steps and 25 Transformer-assisted steps, resulting in a 2.8$\times$ speedup. This aligns closely with the observed time savings illustrated in Figure 4 (upper). In the quadrotor control scenario, for a simulation with 10,000 total simulation steps and 500 MPC control computations, our method substantially reduces the computation time from 237 s (pure iLQR computation at every MPC step) to 13.3 s by combining 1 step of iLQR computation with 49 subsequent steps predicted by the Transformer. This leads to a 17.8× speedup. Additionally, when compared to state-of-the-art OCP solvers OSQP, ECOS, and SCS reported in [44], our approach achieves speedups of 2.49×, 3.63×, and 2.4×, respectively.
@@
Table II shows that our accelerator delivers the expected performance in Transformer inference. In the cart-pole control scenario, it achieves similar speed to the Apple M4 Pro@3.5 GHz, while running 1.55$\times$ faster than the NVIDIA RTX4070 and 17.67$\times$ faster than the Cortex-A72 in the Raspberry Pi. In the quadrotor control scenario, our accelerator is 1.8$\times$ faster than the M4 Pro CPU and 20.8$\times$ faster than the Raspberry Pi. Although the inference time is close to that of the RTX4070 in the quadrotor task, our accelerator uses much less power, consuming only 1.15 W for the cart-pole and 1.68 W for the quadrotor, compared to the steady 13 W drawn by the GPU in both cases.
The hardware resource overhead is extremely low on a mid-range FPGA, with average utilization remaining below 10% for all resource types except BRAM, which is primarily used to store model parameters. Under the same 20nm process, the estimated silicon area is significantly smaller than that of typical CPUs or GPUs.
TABLE II: FPGA Accelerator Performance Metrics
@@
Latency (ms) 1.05 1.73
Speed-Up (vs. 10-core M4 Pro CPU) 0.99$\times$ 1.8$\times$
Speed-Up (vs. RTX4070 GPU) 1.55$\times$ 1.01$\times$
Speed-Up (vs. Quad-core Cortex-A72) 17.67$\times$ 20.8$\times$
Power Reduction (vs. RTX4070 GPU) 11.3$\times$ 7.74$\times$
@@
BRAM Usage (%) 24 41
DSP Usage (%) 4 9
FF Usage (%) 4 3
LUT Usage (%) 9 8
@@ V CONCLUSION
In this paper, we presented Quattro, a Transformer-accelerated iLQR framework that addresses the sequential bottleneck in traditional iLQR algorithms through parallel inference of control gains. By predicting intermediate computational variables, Quattro maintains the structure of optimal control while significantly reducing computation time. Our experiments on cart-pole and quadrotor systems demonstrate that Quattro achieves up to 27$\times$ per-iteration speedup and 17.8$\times$ end-to-end acceleration within an MPC framework. FPGA-based deployment further shows 17–20$\times$ speedups over edge CPUs and up to 1.55$\times$ than GPU with 11.3$\times$ power reduction and dramatically low hardware overhead. While larger Transformer models can yield even higher prediction accuracy (sometimes surpassing the original iLQR), they introduce higher latency and hardware overhead. Therefore, we select a lightweight configuration (dmodel=128, nhead=4) that balances accuracy, efficiency, and employability. These results confirm the potential of Transformer-based acceleration for real-time optimal control. Future work includes extending Quattro to more complex robotic systems, investigating adaptive prompt tuning, and exploring integration with learning-based control for greater robustness in dynamic environments.